OpenStudy (teller):

Would someone care to help me ? Suppose that there are two types of tickets to a show: advance and same-day. The combined cost of one advance ticket and one same-day ticket is $65. For one performance, 25 advance tickets and 40 same-day tickets were sold. The total amount paid for the tickets was $2150. What was the price of each kind of ticket?

OpenStudy (arindameducationusc):

easy.... x+y=65 25x+40y=2150. solve the equations

OpenStudy (arindameducationusc):

you there?

OpenStudy (teller):

30 and 35?

OpenStudy (arindameducationusc):

yes correct

OpenStudy (teller):

thank you!

OpenStudy (teller):

@arindameducationusc mind helping me with another one?

OpenStudy (arindameducationusc):

yes.. i will try

OpenStudy (teller):

A family has two cars. The first car has a fuel efficiency of 40 miles per gallon of gas and the second has a fuel efficiency of 20 miles per gallon of gas. During one particular week, the two cars went a combined total of 1300 miles, for a total gas consumption of 40 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (teller):

now would I set it up like this 40x+20 y and then 1300x+40y?

OpenStudy (teller):

@arindameducationusc

OpenStudy (arindameducationusc):

hey, i m there.... just now

OpenStudy (teller):

Did i set it up right?

OpenStudy (teller):

?

OpenStudy (arindameducationusc):

analysing.....

OpenStudy (teller):

ok

OpenStudy (arindameducationusc):

one thing is sure, your equations are not right... but the correct equations i am wondering.....

OpenStudy (arindameducationusc):

just wait

OpenStudy (teller):

ok

OpenStudy (teller):

I don't have that long...

OpenStudy (arindameducationusc):

got it

OpenStudy (arindameducationusc):

you there?

OpenStudy (arindameducationusc):

sorry for taking time

OpenStudy (teller):

yes im here

OpenStudy (teller):

so what is it? @arindameducationusc

OpenStudy (arindameducationusc):

40x+20y=1300 x+y=40

OpenStudy (teller):

25 and 15??

OpenStudy (arindameducationusc):

yes

OpenStudy (arindameducationusc):

If you liked it don't forget to medal and fan, Thank you...